OpenStudy (anonymous):

evaluate if a=2 and b=-5 (3a)^2

OpenStudy (anonymous):

36

OpenStudy (aroub):

explaination : 3*2 (u subsitute the 2 by a ) 6 squared is equal to 36

OpenStudy (aroub):

explaination : 3*2 (u subsitute the 2 by a ) 6 squared is equal to 36

OpenStudy (anonymous):

Fail for me. Lol 3 * 2 = 6 6 ^ 2 = 6 * 6 36.